OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±3) and foci at (0, ±7)

OpenStudy (tkhunny):

It will look like this: \(\dfrac{(x-h)^{2}}{a^{2}}-\dfrac{(y-k)^{2}}{b^{2}} = 1\) A quick look at the vertices and foci suggests h = k = 0 Now it looks like this \(\dfrac{x^{2}}{a^{2}}-\dfrac{y^{2}}{b^{2}} = 1\) Another quick look at the vertices give a = 3 Now, it looks like this: \(\dfrac{x^{2}}{3^{2}}-\dfrac{y^{2}}{b^{2}} = 1\) Of course, this is not "Standard" form, but there is only on question left to answer. How can the vertices and foci help us fin 'b'?

OpenStudy (anonymous):

so would it be y^2/49 - x^2/9

OpenStudy (tkhunny):

1) You changed the orientation of the hyperbola. That's no good. 2) You used the foci distance for the co-vertex distance. That's no good. 3) It doesn't equal anything. Thus, it's not an equation or relation, just a statement, Let's get back to where I left off. No more guessing. 4) You should have a relationship between a, b, and c for an hyperbola. What is it.